# ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection

### v2

@@ A.2 Implementation details
The computationally most expensive operation in the neural projection layer is the matrix inversion $(BB^{T})^{-1}$ (Def. 1 in Section 4.1), which has a complexity of $\mathcal{O}(N^{3})$. At inference time, $N=N_{C}$, since $B\in\mathbb{R}^{N_{C}\times N_{O}}$ and $BB^{T}\in\mathbb{R}^{N_{C}\times N_{C}}$. Hence, considering a number of constraint $N_{C}<10^{3}$, the matrix inversion is performed in $\sim 1$ million FLOPs, which is a negligible amount for most modern CPUs and GPUs. During training, assuming the use of batch gradient descent and defining the batch size ($BS$) as the number of data points processed in a single iteration, an equivalent number of matrix inversions must be performed. Thus, the computational cost for a single batch is apparently $\mathcal{O}(BS\times N_{C}^{3})$. To address this, we leverage parallel computing on GPUs by constructing a rank-3 tensor $\mathbf{B}\in\mathbb{R}^{BS\times N_{C}\times N_{O}}$ to hold $BS$ local matrices $B$. Similarly, a rank-2 tensor $V\in\mathbb{R}^{BS\times N_{C}}$ is built to store $BS$ local vectors $\mathbf{v}$. Modern deep learning libraries enable batch operations, such as matrix inversion, which reduce the effective complexity to $\mathcal{O}(N_{C}^{3})$ (i.e., scaling only with the number of constraints).
Given these conditions and the capabilities of current hardware, the neural projection operation remains computationally efficient even during training when the number of constraints is fewer than a few thousand. This facilitates new computational approaches for handling and enforcing nonlinear constraints in neural networks.

### v3

@@ D.3 Batch local projection
The computationally most expensive operation in the neural projection layer is the matrix inversion $(BB^{T})^{-1}$ (Def. 1 in Section 4.1), which has a complexity of $\mathcal{O}(N^{3})$. At inference time, $N=N_{C}$, since $B\in\mathbb{R}^{N_{C}\times N_{O}}$ and $BB^{T}\in\mathbb{R}^{N_{C}\times N_{C}}$. Hence, considering a number of constraints $N_{C}<10^{3}$, the matrix inversion is performed in less than 1 million FLOPs, which is an affordable amount for most modern CPUs and GPUs. During training, assuming the use of batch gradient descent and defining the batch size ($BS$) as the number of data points processed in a single iteration, an equivalent number of matrix inversions must be performed. Thus, the computational cost for a single batch is apparently $\mathcal{O}(BS\times N_{C}^{3})$. To address this, we leverage parallel computing on GPUs by constructing a rank-3 tensor $\mathbf{B}\in\mathbb{R}^{BS\times N_{C}\times N_{O}}$ to hold $BS$ local matrices $B$. Similarly, a rank-2 tensor $V\in\mathbb{R}^{BS\times N_{C}}$ is built to store $BS$ local vectors $\mathbf{v}$. Modern deep learning libraries enable batch operations, such as matrix inversion, which reduce the effective complexity to $\mathcal{O}(N_{C}^{3})$ (i.e., scaling only with the number of constraints). To invert the batch of matrices, we use the Cholesky factorization algorithm [39].
Given these conditions and the capabilities of current hardware, the neural projection operation remains computationally efficient even during training when the number of constraints is in the order of a few hundred. Moreover, the complexity of this method is equivalent to other state-of-the-art methods such as DC3 [7], where each Newton’s step in the completion algorithm requires the inversion of a batch of $(N_{C}\times N_{C})$ matrices.
D.4 Memory footprint
For each neural projection layer, the AdaNP module creates four 3D tensors:
  • Tensor $\mathbf{B}$ of shape $(BS,N_{C},N_{O})$
  • Tensor $\mathbf{b}$ of shape $(BS,N_{C},1)$
  • Tensor $\mathbf{B^{*}}$ of shape $(BS,N_{O},N_{O})$
  • Tensor $\mathbf{b^{*}}$ of shape $(BS,N_{O},1)$
Assuming 32-bit floating-point representation (i.e., 4 bytes per element), the memory required to store these tensors for a single projection step is:
[…]
Figures 6 and 7 illustrate how the memory requirement for a single projection layer is affected by variations in batch size, the number of predicted variables, and the number of constraints.
The total memory usage of the unrolled AdaNP depends on the mode of operation:
  • During training, tensors are retained at each of the $n$ projection steps (i.e., for gradient computation), resulting in:
    $M_{\text{train}}=4\cdot n\cdot BS\cdot(N_{O}+1)\cdot(N_{C}+N_{O})$ (73)
  • During inference, tensors are not retained across steps, and the peak memory usage is:
    $M_{\text{infer}}=4\cdot BS\cdot(N_{O}+1)\cdot N_{O}$ (74)
Figure 6: Memory footprint of a single projection layer $\mathcal{P}$. $N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output. The memory usage scales linearly with the batch size, with the growth rate determined by the number of predicted variables and constraints. For a batch size of 200, the memory requirement is on the order of tens of megabytes.
Figure 7: Memory footprint for a fixed batch size of 200 samples. $N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output. For large-scale tasks involving thousands of variables and constraints, the AdaNP module can become memory-intensive, with each projection layer requiring more than 1 GB of memory.
Figure 8: Memory usage during training and inference as a function of the number of projection layers $n$, for two problem scales. $N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output. Results are shown for a fixed batch size $BS=200$. Smaller-scale setup: $N_{C}=50$, $N_{O}=100$; larger-scale setup: $N_{C}=150$, $N_{O}=200$. Training memory increases linearly with $n$, with significantly higher usage in larger-scale problems. Inference memory remains nearly constant across all configurations.
Appendix E Outlook and limitations
The method has the potential to address specific challenges or complement existing approaches, including those based on NNs and other machine learning models (e.g., Gaussian processes and support vector machines). For instance, hard constraints can be combined with soft-constraint techniques, such as PINNs, to reliably solve PDEs [62]. Additionally, the method could enhance learning performance in partially annotated datasets by inferring missing information through available constraints. Finally, an interesting future direction could involve applying the AdaNP module to GenAI models, guiding the generation process toward domain-compliant samples, such as for synthetic data or image/video generation.
Limitations
In this paragraph, we highlight the main limitation of the proposed method. Firstly, the effectiveness of ENFORCE is highly dependent on the regression capabilities of the neural network backbone. When the model lacks sufficient complexity to achieve accurate predictions, ENFORCE provides limited benefit. This is supported by the theoretical implications of the orthogonal projections on the constraints manifold discussed in this study. Furthermore, the method becomes computationally and memory-intensive when applied to systems with a large number of constraints (e.g., more than a few hundred). This is due to the computational cost of each neural projection layer, which scales as $O(N_{C}^{3})$ because of the matrix inversion operation, where $N_{C}$ is the number of constraints. Further details are provided in Appendix D.